# Joint Source-Channel Optimization for UAV Video Coding and Transmission

### v1

@@ II-B2 Coding Delay and Power Models
In the research field of video coding, encoding complexity is an important indicator of quantization encoding delay. Compared to AVC, HEVC employs flexible coding units (CU) partitioning based on a quadtree structure and a richer set of coding modes. It results in relatively higher computational complexity for CU recursive detection and coding mode selection. However, researchers have conducted extensive research on the CU partitioning and the rapid selection of coding modes. As a result, the computational load of recursive detection and coding mode selection has been significantly reduced[33, 34, 35, 36].
In the block-based hybrid video coding framework, ME has the highest computational complexity and is dominant throughout the entire encoding process. Thus, the computational complexity of ME is utilized to approximate the complexity of the entire video coding process. Further, in HEVC, to achieve higher compression efficiency, more complex inter-frame prediction modes and tree-structured motion compensation mechanisms are adopted. Compared with AVC, the time consumption of ME in HEVC accounts for a larger proportion. Therefore, this paper employs motion estimation time (MET) to approximate the encoding latency of HEVC.
@@ II-C1 Rate and Distortion Models
In the UAV deployment scenario considered in this paper, signals received by the ECV are predominantly LoS signals. Then, the bit error rate (BER) curve of signals received by the ECV approximates to that of an additive white Gaussian noise (AWGN) channel. Consequently, the BER model of the AWGN channel is employed to approximate the BER of signals received by the ECV[38], which can be expressed as below
@@ II-C2 Transmission Delay and Power Models
Transmission delay ${t_{trans}}$ refers to the total time it takes for a data packet to travel from the sender of the channel, through the transmission process, and ultimately reach the receiver. It mainly includes sending delay ${t_{send}}$, propagation delay ${t_{propa}}$, queuing delay ${t_{queue}}$, and buffer processing delay ${t_{buffer}}$, as illustrated in Fig. 2.
@@
Propagation delay ${t_{propa}}$ is the time required for an electromagnetic signal to travel a certain distance within a communication channel. The calculation formula is ${t_{propa}}=\frac{d}{c}$, where represents the signal propagation distance between ECV and UAV, and denotes the propagation speed of the electromagnetic wave in the medium. In the application scenario discussed in this paper, since the propagation distance between the UAV and the ECV is relatively short, the propagation delay can be neglected.
Queuing delay ${t_{queue}}$ is the time a data packet spends waiting to be processed at network nodes such as switches or routers during network transmission. Specifically, upon arrival at a switch or router, a data packet must first wait in the input queue for processing. After the switch or router determines the forwarding interface, the packet also needs to wait in the output queue for forwarding. In the application scenario considered in this paper, since there is no forwarding operation involving switches or routers, it can be considered that there is no queuing delay.
@@ III-A Problem Decomposition
(17) indicates that the minimization of the drift-plus-penalty term can be approximated by minimizing its upper bound. Consequently, the complex sequential-decision problem can be approximately transformed into a family of repeated optimization problems aiming at minimizing the upper bound of the drift-plus-penalty function. Specifically, we can repeatedly optimize the following problem at each time slot $t$ to obtain the upper bound of the objective function of (18).
@@ III-B Solution to the Quantization Step Sub-Problem
After substituting the R-D model into (19), it can be observed that the expression of the objective function is complex, making it difficult to analyze and optimize directly. To facilitate the analysis of the objective function of this sub-problem, this paper introduces a slack variable $\Omega$ and sets ${R_{e}}(\lambda,Q;t)\leq\Omega$. Meanwhile, a slack variable $\delta$ is introduced, and set $\delta\geq C\cdot{({R_{e}}(\lambda,Q;t))^{-K}}$. Consequently, (19) can be reformulated as
@@
In summary, this paper introduces a set of slack variables $\delta$, $\varepsilon$, $\xi$, and $\Omega$, and employs the SCA strategy to perform convex approximations on the non-convex constraints. As a result, we can approximately transform (19) into the following convex optimization problem.
@@ III-C Solution to the Power Control and Search Range Sub-Problem
In summary, this paper introduces a series of slack variables $\delta$, $\varepsilon$, $\xi$, $\Omega$, $\varphi$, $\tau$, and $\zeta$ to perform relaxation transformations on the optimization problem, and employs the SCA strategy to perform convex approximations on the non-convex constraints. (24) can then be transformed into the following convex optimization problem
@@ III-D Algorithm Design
1:  Initialization: Initialize a positive random value for $X(1)$, set the maximum UAV transmit power ${P_{\max}}$, the maximum tolerable delay ${d_{\max}}$, the maximum number of time slots $T$, and the maximum number of iterations ${r_{\max}}$.
@@
At each time slot, the computational complexity of Algorithm 1 is primarily contributed by two components: the optimization sub-problem of the quantization step and the optimization sub-problem of transmit power and search range. These two sub-problems are transformed into convex problems and solved using an interior method, with their respective computational complexity being $O({N^{3.5}})$ and $O({M^{3.5}})$, where $N$ and $M$ are the dimensions of the decision variables of sub-problems. Additionally, the iterative optimization process of these two sub-problems needs to be considered. In the worst case, the total computational complexity of Algorithm 1 is $O({r_{\max}}(O({N^{3.5}})+O({M^{3.5}})))$, where ${r_{\max}}$ represents the maximum number of iterations.
@@ IV-A Experimental Environment and Parameter Setting
In this section, we will validate the effectiveness of the proposed algorithm. The experiments utilize the official standard test model of HEVC, namely the latest version of HM, HM16.20. The experiments select standard video test sequences City (CIF, $352\times 288$) and Coastguard (CIF, $352\times 288$) for testing. The City sequence captures stationary targets by moving the camera position, while the Coastguard sequence tracks and shoots moving targets with a fixed camera. Both test sequences can effectively simulate typical application scenarios of UAV video surveillance.
In this paper, obtaining the regression coefficients through experimentation is a critical step. Initially, the transformation residual data of video sequences is acquired. The experiments utilize the low-latency configuration file encoder_lowdelay_P_main.cfg of HM, with the search range $\lambda\in\{1,4,8,16,32\}$ and the quantization parameter $QP\in\{18,24,30,36,42\}$. The encoding is performed on the basis of different $\lambda$ and $QP$. During the encoding process, the transformed residual values of each prediction unit in P frames are written to files. A total of $25$ transformed residual data files can be obtained, corresponding to each $(\lambda,QP)$ pair. Subsequently, for each transformed residual data file, the standard deviation of residuals for all prediction units is calculated on a per-frame basis. To better capture video characteristics and mitigate the adverse effects on the experimental results due to fluctuations in inter-frame transformation residual standard deviations, the average of the transformed residual standard deviations for the first $10$ frames of the video sequence is calculated. Through the above calculations, experimental values of the residual standard deviation corresponding to different combinations of $\lambda$ and $QP$ can be obtained. Finally, based on the previously calculated residual standard deviations, nonlinear regression analysis is performed to determine values of coefficients such as ${a_{1}}$, ${a_{2}}$, ${a_{3}}$, and ${a_{4}}$ in (3), thereby obtaining the closed-form expression for $\sigma(\lambda,Q)$.
The configuration of other parameters for the experiments is as follows: $V=4$, ${\rho_{1}}=2$, ${\rho_{2}}=0.01$, $C=0.0015$, $K=0.55$, $T=40$, $\mu=0.1$, maximum total power consumption ${P_{\max}}=2000$ mW, data block length $L=1$ Mb, clock frequency ${F_{clk}}=1000$ MHz, frame rate of $30$ fps, network bandwidth $B=10$ MHz, the position of the ECV $\left[{50\ \rm{m},50\ \rm{m}}\right]$, the UAV flight radius $250$ m, the UAV trajectory center position $\left[{250\ \rm{m},250\ \rm{m},500\ \rm{m}}\right]$, and the UAV flight line speed $20$ m/s. Other parameters related to the AtG channel model can be referred to[40].
@@ IV-B Performance Evaluation
In this section, we design extensive experiments to validate the performance of the proposed algorithm. These experiments include verification of the algorithm’s stability, analysis of variations in the search range $\lambda$ and UAV transmit power ${P_{t}}(t)$, assessment of the d-P-R-D model, comparative analysis of the optimization results of the proposed algorithm, and a comparative analysis of the performance of the proposed algorithm with HM rate control (HM RC) algorithm.
@@ IV-B1 Stability of LyaRI
Figure 4: Convergence of $\lambda$ and $Q$ as well as the stability of $X(t)$ vs. time slots.
Fig. 4 illustrates the convergence behavior of LyaRI on the Coastguard sequence in terms of search range $\lambda$ and quantization step $Q$, as well as the stability of virtual queue $X(t)$. Here, we set ${d_{\max}}=3$s, and set two combinations of different initial values to be ${\left({\lambda,Q,{P_{t}}}\right)_{\min}}=(4,30,100\ \rm{mW})$ and ${\left({\lambda,Q,{P_{t}}}\right)_{\max}}=(16,140,1200\ \rm{mW})$, respectively. Experimental results demonstrate that both $\lambda$ and $Q$ can rapidly converge to their optimal values after a finite number of iterations. For different initial value combinations of $\lambda$, $Q$ and ${P_{t}}(t)$, although the number of iterations varies, the parameter ultimately converges to the same optimal value.
@@
IV-B2 Joint Optimization Performance Analysis
@@
We can observe from this figure that: Firstly, $\lambda$ and ${P_{t}}(t)$ exhibit periodic changes and show an approximately symmetrical characteristic. This is because the distance between the UAV and the ECV changes periodically when the UAV completes a full circular trajectory flight.
@@
Figure 6: Source distortion vs. ${d_{\max\_trans}}$ curves for different ${d_{\max}}$.
@@
(14c) shows that the tolerable transmission delay ${d_{\max\_trans}}$ is inversely proportional to the UAV data rate. Therefore, this experiment adopts ${d_{\max\_trans}}$ to map the data rate. To verify the interrelationships among the elements in d-P-R-D model under the condition of a given maximum power ${P_{\max}}$, this experiment obtains the source distortion (SD) curves with respect to ${d_{\max\_trans}}$ for different maximum delays ${d_{\max}}$.
@@
For Coastguard sequence, we set ${d_{\max}}\in\left\{{2.75,2.80,2.85,2.90}\right\}$. We choose a value for ${d_{\max}}$, vary the value of ${d_{\max\_trans}}$, and obtain the corresponding value of $SD$, thus acquiring the $SD-{d_{\max\_trans}}$ curve corresponding to ${d_{\max}}$. Fig. 6(a) illustrates the $SD-{d_{\max\_trans}}$ curves for Coastguard sequence under different ${d_{\max}}$ settings. The observation results show that, for the same ${d_{\max}}$, the larger the ${d_{\max\_trans}}$, the greater the $SD$. This is because a larger ${d_{\max\_trans}}$ implies a relatively smaller data rate ${R_{c}}(t)$. According to (14b), the source coding rate ${R_{e}}(\lambda,Q;t)$ is also relatively smaller, thus leading to a larger $SD$. Besides, it can be observed that for the same ${d_{\max\_trans}}$, the larger the ${d_{\max}}$, the smaller the $SD$. This is because, for the same ${d_{\max\_trans}}$, a larger ${d_{\max}}$ implies a greater source encoding delay and a smaller $SD$. Experimental results are consistent with the theoretical analysis of the video coding d-P-R-D model. It validates the effectiveness of the designed model.
For City sequence, we set ${d_{\max}}\in\left\{{2.10,2.15,2.20,2.25}\right\}$. Fig. 6(b) demonstrates that the trend of City sequence’s curve is consistent with that of Coastguard sequence; hence, no detailed analysis is presented here.
@@
IV-B4 Comparative Analysis of Optimization Results of LyaRI
This experiment is primarily conducted to validate the encoding performance of LyaRI. Initially, the optimized video coding parameter pair $({\lambda^{*}},Q{P^{*}})$ is obtained by LyaRI. Subsequently, based on $({\lambda^{*}},Q{P^{*}})$, new encoding parameter pairs are configured by incrementing or decrementing their values. Thereafter, a comparative verification is conducted within HM by assessing Y-PSNR, coding bitrate, and encoding time for each parameter pair during the first $10$ time slots.
For Coastguard sequence, we set $d\max=3$s, and the optimized parameter pair obtained by LyaRI is $({\lambda^{*}},Q{P^{*}})=(9,41)$. With the optimized value ${\lambda^{*}}=9$ being fixed, we change the value of $QP$. For instance, setting $(\lambda,QP)=(9,40)$, it can be observed from Fig. 7 that Y-PSNR is higher than the optimized solution. However, this type of parameter configuration violates the bitrate constraint. It indicates that the parameter configuration $(\lambda,QP)=(9,40)$ is infeasible. On the other hand, when $QP$ is fixed at the optimized value of $41$, and $\lambda$ is increased, such as set $\lambda=16$. In this case, Y-PSNR and bitrate performance do not show significant improvement compared to the optimized solution. Nevertheless, the encoding time is much greater than that of the optimized solution.
The search range $\lambda$ significantly affects the complexity of video coding. After obtaining the optimized value of $\lambda$ using LyaRI, it eliminates the need to allocate valuable computational resources to additional ME. This is particularly beneficial in saving encoding time and reducing power consumption. Additionally, setting a lower $\lambda$, such as $\lambda=1$, results in a Y-PSNR that is lower than the optimized solution, leading to a degradation in encoding performance. For City sequence, we set $d\max=2$s, and the optimized parameter pair obtained by LyaRI is $({\lambda^{*}},Q{P^{*}})=(7,40)$. From Fig. 7, it can be observed that the encoding performance trend of City sequence is consistent with that of Coastguard sequence; hence, no further analysis is provided here.

### v2

@@ II-B2 Coding Delay and Power Models
In the research field of video coding, encoding complexity is an important indicator of quantization encoding delay. Compared to AVC, HEVC employs flexible coding units (CU) partitioning based on a quadtree structure and a richer set of coding modes. It results in relatively higher computational complexity for CU recursive detection and coding mode selection. However, the CU partitioning and the rapid selection of coding modes have been extensively researched. As a result, the computational load of recursive detection and coding mode selection has been significantly reduced[33, 34, 35, 36].
In the block-based hybrid video coding framework, ME has the highest computational complexity and is dominant throughout the entire encoding process. Thus, the computational complexity of ME is utilized to approximate the complexity of the entire video coding process [16]. Further, in HEVC, to achieve higher compression efficiency, more complex inter-frame prediction modes and tree-structured motion compensation mechanisms are adopted compared with AVC. Therefore, this paper also employs motion estimation time (MET) to approximate the encoding latency of HEVC.
@@ II-C1 Rate and Distortion Models
In the UAV deployment scenario considered in this paper, signals received by the ECV are predominantly LoS signals. Then, the bit error rate (BER) curve of signals received by the ECV approximates to that of an additive white Gaussian noise (AWGN) channel. Thus, the BER model of the AWGN channel is employed to approximate the BER of signals received by the ECV[38], which can be expressed as below
@@ II-C2 Transmission Delay and Power Models
Transmission delay ${t_{trans}}$ refers to the total time it takes for a data packet to travel from the sender of the channel, through the transmission process, and ultimately reach the receiver. It mainly includes sending delay ${t_{send}}$, propagation delay ${t_{propa}}$, and buffer processing delay ${t_{buffer}}$, as illustrated in Fig. 2.
@@
Propagation delay ${t_{propa}}$ is the time required for an electromagnetic signal to travel a certain distance within a communication channel. The calculation formula is ${t_{propa}}=\frac{d}{c}$, where $d$ represents the signal propagation distance between the UAV and the ECV, and $c$ denotes the propagation speed of the electromagnetic wave in the medium. In the application scenario discussed in this paper, since the propagation distance between the UAV and the ECV is relatively short, the propagation delay can be neglected.
@@ III-A Problem Decomposition
(17) indicates that the minimization of the drift-plus-penalty term can be approximated by minimizing its upper bound. As a result, the complex sequential-decision problem can be approximately transformed into a family of repeated optimization problems aiming at minimizing the upper bound of the drift-plus-penalty function. Specifically, we can repeatedly optimize the following problem at each time slot $t$ to obtain the upper bound of the objective function of (18).
@@ III-B Solution to the Quantization Step Sub-Problem
After substituting the R-D model into (19), it can be observed that the expression of the objective function is complex, making it difficult to analyze and optimize directly. To facilitate the analysis of the objective function of this sub-problem, this paper introduces a slack variable $\Omega$ and sets ${R_{e}}(\lambda,Q;t)\leq\Omega$. Meanwhile, a slack variable $\delta$ is introduced, and we set $C\cdot{({R_{e}}(\lambda,Q;t))^{-K}}\leqslant\delta$. Thus, (19) can be reformulated as
@@
In summary, this paper introduces a set of slack variables $\Omega$, $\delta$, $\varepsilon$, and $\xi$, and employs the SCA strategy to perform convex approximations on the non-convex constraints. As a result, we can approximately transform (19) into the following convex optimization problem.
@@ III-C Solution to the Power Control and Search Range Sub-Problem
In summary, this paper introduces a series of slack variables $\Omega$, $\delta$, $\zeta$, $\varepsilon$, $\xi$, $\varphi$, and $\tau$ to perform relaxation transformations on the optimization problem, and employs the SCA strategy to perform convex approximations on the non-convex constraints. (24) can then be transformed into the following convex optimization problem
@@ III-D Algorithm Design
1:  Initialization: Initialize a positive random value for $X(1)$, set the maximum UAV total power ${P_{\max}}$, the maximum tolerable delay ${d_{\max}}$, the maximum number of time slots $T$, and the maximum number of iterations ${r_{\max}}$.
@@
At each time slot, the computational complexity of Algorithm 1 is primarily contributed by two components: the optimization of the quantization step sub-problem and the optimization of transmit power and search range sub-problem. These two sub-problems are transformed into convex problems and solved using an interior method, with their respective computational complexity being $O({N^{3.5}})$ and $O({M^{3.5}})$, where $N$ and $M$ are the dimensions of the decision variables of sub-problems. Additionally, the iterative optimization process of these two sub-problems needs to be considered. In the worst case, the total computational complexity of Algorithm 1 is $O({r_{\max}}(O({N^{3.5}})+O({M^{3.5}})))$, where ${r_{\max}}$ represents the maximum number of iterations.
@@ IV-A Experimental Environment and Parameter Setting
In this section, we will validate the effectiveness of the proposed algorithm. The experiments utilize the latest version of official standard test model of HEVC (HM), namely HM 16.20. The experiments select standard video sequences City (CIF, $352\times 288$) and Coastguard (CIF, $352\times 288$) for testing. The City sequence captures stationary targets by moving the camera position, while the Coastguard sequence tracks and shoots moving targets with a fixed camera. Both test sequences can effectively simulate typical UAV video surveillance scenarios.
In this paper, obtaining the regression coefficients through experimentation is a critical step. Initially, the transformed residual data of video sequences is acquired. The experiments utilize the low-latency configuration file encoder_lowdelay_P_main.cfg of HM, with search range $\lambda\in\{1,4,8,16,32\}$ and quantization parameter $QP\in\{18,24,30,36,42\}$. The encoding is performed on the basis of different $\lambda$ and $QP$. During the encoding process, the transformed residual values of each prediction unit in P frames are written to files. A total of $25$ transformed residual data files can be obtained, corresponding to each $(\lambda,QP)$ pair. Subsequently, for each transformed residual data file, the standard deviation of residuals for all prediction units is calculated on a per-frame basis. To better capture video characteristics and mitigate the adverse effects on the experimental results due to fluctuations in inter-frame transformed residual standard deviations, the average of the transformed residual standard deviations for the first $10$ frames of the video sequence is calculated. Through the above calculations, experimental values of the residual standard deviation corresponding to different combinations of $\lambda$ and $QP$ can be obtained. Finally, based on the previously calculated residual standard deviations, nonlinear regression analysis is performed to determine values of coefficients such as ${a_{1}}$, ${a_{2}}$, ${a_{3}}$, and ${a_{4}}$ in (3), thereby obtaining the closed-form expression for $\sigma(\lambda,Q)$.
The configuration of other parameters for the experiments is as follows: $V=4$, ${\rho_{1}}=2$, ${\rho_{2}}=0.01$, $C=0.0015$, $K=0.55$, $T=40$, $\mu=0.1$, maximum total power consumption ${P_{\max}}=2000$ mW, data block length $L=1$ Mb, encoding power-related constant $k=1.3\times 10^{-24}$ mW/Hz, clock frequency ${F_{clk}}=1000$ MHz, and network bandwidth $B=10$ MHz. We let the frame rate be $30$ fps, the position of the ECV be $\left[{50\ \rm{m},50\ \rm{m}}\right]$, the UAV flight radius be $250$ m, the UAV trajectory center position be $\left[{250\ \rm{m},250\ \rm{m},500\ \rm{m}}\right]$, and the UAV flying speed be $20$ m/s. Other parameters related to the AtG channel model can be referred to[40].
@@ IV-B Performance Evaluation
In this section, we design extensive experiments to validate the performance of LyaRI. These experiments include verification of the stability of the algorithm, joint optimization performance verification, an assessment of the d-P-R-D model, a comparative analysis of the optimization results of LyaRI, a comparative analysis of the performance of LyaRI and HM rate control (HM RC) algorithm, and a comparative analysis of subjective performance.
@@ IV-B1 Stability of LyaRI
Figure 4: Convergence performance of $\lambda$ and $Q$, and the stability of $X(t)$.
Fig. 4 illustrates the convergence behavior of LyaRI on Coastguard sequence in terms of search range $\lambda$ and quantization step $Q$, as well as the stability of virtual queue $X(t)$. Here, we set ${d_{\max}}=3$s, and set two combinations of different initial values to be ${\left({\lambda,Q,{P_{t}}}\right)_{\min}}=(4,30,100\ \rm{mW})$ and ${\left({\lambda,Q,{P_{t}}}\right)_{\max}}=(16,140,1200\ \rm{mW})$, respectively. Experimental results demonstrate that both $\lambda$ and $Q$ can rapidly converge to their optimal values after a finite number of iterations. For different initial value combinations of $\lambda$, $Q$ and ${P_{t}}(t)$, although the number of iterations varies, the parameter ultimately converges to the same optimal value.
@@
IV-B2 Joint Optimization Performance Verification
@@
From Fig. 5, we can observe that: Firstly, $\lambda$ and ${P_{t}}(t)$ exhibit periodic changes and show an approximately symmetrical characteristic. This is because the distance between the UAV and the ECV changes periodically when the UAV completes a full circular trajectory flight.
@@
Figure 6: Source distortion vs. ${d_{\max\_trans}}$ curves with different ${d_{\max}}$.
@@
(14c) shows that the tolerable transmission delay ${d_{\max\_trans}}$ is inversely proportional to the UAV data rate. Therefore, this experiment adopts ${d_{\max\_trans}}$ to map the data rate. To verify the interrelationships among the elements in d-P-R-D model under the condition of a given maximum total power ${P_{\max}}$, this experiment obtains the source distortion (SD) curves with respect to ${d_{\max\_trans}}$ for different maximum delays ${d_{\max}}$.
(a) Trends in Y-PSNR, varying QP and $\lambda$
(b) Trends in bitrate, varying QP and $\lambda$
(c) Trends in encode time, varying QP and $\lambda$
(d) Trends in Y-PSNR, varying QP and $\lambda$
(e) Trends in bitrate, varying QP and $\lambda$
(f) Trends in encode time, varying QP and $\lambda$
@@
For Coastguard sequence, we set ${d_{\max}}\in\left\{{2.75,2.80,2.85,2.90}\right\}$ s. We choose a value for ${d_{\max}}$, vary the value of ${d_{\max\_trans}}$, and obtain the corresponding value of $SD$, thus acquiring the $SD-{d_{\max\_trans}}$ curve corresponding to ${d_{\max}}$. Fig. 6(a) illustrates the $SD-{d_{\max\_trans}}$ curves for Coastguard sequence under different ${d_{\max}}$ settings. The observation results show that, for the same ${d_{\max}}$, the larger the ${d_{\max\_trans}}$, the greater the $SD$. This is because a larger ${d_{\max\_trans}}$ implies a relatively smaller data rate ${R_{c}}(t)$. According to (14b), the source coding rate ${R_{e}}(\lambda,Q;t)$ is also relatively smaller, thus leading to a larger $SD$. Besides, it can be observed that for the same ${d_{\max\_trans}}$, the larger the ${d_{\max}}$, the smaller the $SD$. This is because, for the same ${d_{\max\_trans}}$, a larger ${d_{\max}}$ implies a greater source encoding delay, and thus a smaller $SD$. Experimental results are consistent with the theoretical analysis of the video coding d-P-R-D model. It validates the effectiveness of the designed model.
For City sequence, we set ${d_{\max}}\in\left\{{2.10,2.15,2.20,2.25}\right\}$ s. Fig. 6(b) demonstrates that the trend of City sequence’s curve is consistent with that of Coastguard sequence; hence, no detailed analysis is presented here.
(a) Trends in Y-PSNR, varying initial QP
(b) Trends in bit, varying initial QP
(c) Trends in encode time, varying initial QP
(d) Trends in Y-PSNR, varying initial QP
(e) Trends in bit, varying initial QP
(f) Trends in encode time, varying initial QP
@@
IV-B4 Comparative Analysis of LyaRI Optimization Results
This experiment is primarily conducted to validate the encoding performance of LyaRI. Initially, the optimized video coding parameter pair $({\lambda^{*}},Q{P^{*}})$ is obtained by LyaRI. Subsequently, based on $({\lambda^{*}},Q{P^{*}})$, new encoding parameter pairs are configured by incrementing or decrementing their values. Thereafter, a comparative verification is conducted within HM by assessing Y-PSNR, encoding bitrate, and encoding time for each parameter pair during the first $10$ time slots.
For Coastguard sequence, we set $d_{\max}=3$ s, and the optimized parameter pair obtained by LyaRI is $({\lambda^{*}},Q{P^{*}})=(9,41)$. With the optimized value ${\lambda^{*}}=9$ being fixed, we change the value of $QP$. For instance, setting $(\lambda,QP)=(9,40)$, it can be observed from Fig. 7 that Y-PSNR is higher than the optimized solution. However, this type of parameter configuration violates the bitrate constraint. It indicates that the parameter configuration $(\lambda,QP)=(9,40)$ is infeasible. On the other hand, when $QP$ is fixed at the optimized value of $41$, and $\lambda$ is increased, such as set $\lambda=16$. In this case, Y-PSNR and bitrate performance do not show significant improvement compared to the optimized solution. Nevertheless, the encoding time is much greater than that of the optimized solution.
The search range $\lambda$ significantly affects the complexity of video coding. After obtaining the optimized value of $\lambda$ using LyaRI, it eliminates the need to allocate valuable computational resources to additional ME. This is particularly beneficial in saving encoding time and reducing power consumption. Additionally, setting a lower $\lambda$, such as $\lambda=1$, results in a lower Y-PSNR value relative to the optimized solution. For City sequence, we set $d_{\max}=2$ s, and the optimized parameter pair obtained by LyaRI is $({\lambda^{*}},Q{P^{*}})=(7,40)$. From Fig. 7, it can be observed that the encoding performance trend of City sequence is consistent with that of Coastguard sequence; hence, no further analysis is provided here.